# LSHBloom: Internet-Scale Text Deduplication

### v3

@@ 5. Evaluation
Broadly our evaluation considers two concerns important for large-scale deduplication: correctness and performance. As expressed above, both the correctness and performance of these methods are critical for their application to extremely large datasets. To assess correctness, we tune LSHBloom and other state-of-the-art deduplication methods on a balanced dataset of 25K PDFs containing labeled duplicates. We then use the best parameter settings for each algorithm and compare them on a larger labelled dataset of 50K PDFs. Without sufficient correctness, the system risks either under-identifying duplicates, leading to decreased downstream model quality and unnecessary training overhead; or over-identifying duplicates, which unnecessarily reduces the training dataset volume. For performance, we consider the storage utilization and runtime of each method on a larger collection of 39M PDFs, peS2o (peS2o). Without sufficient performance, it may not be possible to process large datasets on available hardware resources.
In the following subsections, we describe the details of our evaluation beginning with the evaluation methodology, then proceeding to correctness, performance, and finally the tradeoffs between performance and correctness.
@@ 5.1. Experimental Methodology
In this subsection, we describe our experimental methodology including testbed hardware, state-of-the-art methods for comparison, our evaluation metrics, and datasets.
@@ 5.1.2. Baselines
We compare LSHBloom with state-of-the-art deduplication methods widely used in large-scale LLM training: MinHashLSH, Dolma, Dolma-ngram, CCNet, and DataComp-LM. We normalized all implementations to Python-only implementations to avoid performance differences due to implementation language. This also enabled us to normalize the Bloom filter implementation for those techniques which use them. Below, we describe any changes or configurations to the methods.
MinhashLSH (broder1997resemblance) (lsh): We configure the traditional MinhashLSH algorithm as described in subsection 2.3. DOLMA: We apply both standard dolma and dolma-ngram methods to document-level deduplication. Specifically, for dolma we determine whether a document is duplicated by identifying duplicated paragraphs and comparing the percentage of document text duplicated to the overlap threshold, $T$. For dolma-ngram, we split each document into a set of n-grams with a given size, determine how many n-grams are duplicated according to the Bloom filter, and compare the percentage of n-grams that are duplicated within that document to the overlap threshold, $T$. Our implementation aims to faithfully extend Dolma’s two primary deduplication methods to document-level deduplication. CCNet originally performs deduplication at the paragraph level; however, here, we apply it at the document level by measuring the proportion of duplicated paragraphs in any given document against some tolerance threshold, $T\in$ [0, 1]. For a fair comparison, we compare documents to all other documents—just as MinhashLSH does. When evaluating memory usage, we apply all CCNet optimizations, such as the use of memory-efficient hash sets. Additionally, we augment CCNet with multiprocessing where appropriate (e.g., when computing hashes over each document) for a more appropriate comparison during our scaling tests. DataComp-LM: For our evaluation, we compare only with their document-level deduplication procedure.
When considering the correctness of techniques that operate on n-grams and use Bloom filters, we must give an expected n-gram count in the corpus; choosing this count accurately is crucial to make a fair comparison. Because counting the number of n-grams of a given size in a corpus is far more computationally expensive than counting the number of documents (as is required for our technique), we compute an estimate of the n-gram count in a corpus. To do so we randomly sample $N$ = 1000 documents from the corpus, compute the mean n-gram count from our sample, and multiply the mean n-gram count by the number of total documents in that corpus to obtain an estimate of the number of n-grams in the entire corpus. We parameterize Bloom filters for n-gram techniques using this estimate. We apply an analogous procedure to obtain estimated paragraph counts for Dolma and CCNet.
@@ 5.1.3. Metrics
Correctness: We require methods that can distinguish meaningfully between duplicate and non-duplicate documents. Here, we consider a ”positive” label to mean that a document is a duplicate of another document in the corpus. We want these methods to yield a few false positives (to retrain the maximal number of training examples), few false negatives to reduce the prevalence of duplication in our dataset, and to provide a reasonable combination of accuracy and performance. Particularly for large-scale deep learning, retaining the maximal number of training examples is highly desirable—therefore, we want few false positives. Similarly, allowing a large number of duplicates into the training corpus is unacceptable as this affects the model’s training efficiency and may increase the likelihood that our model develops degenerate behaviors such as memorization. To account for these considerations, we evaluate each deduplication technique based on its ability to mitigate both false positives and false negatives using its precision (the proportion of duplicate labels that are true duplicates), its recall (the proportion of duplicates in the dataset that we correctly identify), and its F1 score (the harmonic mean of precision and recall):
@@
Performance: To assess the resource costs of each technique we measure wall clock time and disk space usage. We measure wall clock time using /user/bin/time and assess each method’s disk space requirements using Python’s standard os.path.getsize method.
5.1.4. Constructing the correctness evaluation dataset
To permit the rigorous evaluation of de-duplication methods, we created a pipeline for generating labeled datasets of duplicate documents. We acquire scientific articles (scraped from HTML and parsed from PDF) from a variety of sources, spanning a diverse set of disciplines (mathematics, physics, chemistry, biology, economics, engineering, and medicine) as identified through keywords or classification algorithms based on metadata. Text content from HTML is directly extracted. We translated each PDF into a sequence of tokens by applying a text extraction or OCR tool—one of PyMupDF (PyMuPDF), Nougat (blecher2023nougat), or Tesseract (smith2007overview), each used with roughly the same frequency. We then can construct datasets where documents are duplicated, either by including multiple versions of the same document which were parsed using different text extraction tools and/or by including randomly truncated versions of that document. This reflects common forms of duplication that occur in real-world web scraping pipelines where the same document can be ingested via different text extraction tooling either due to differences in the origin of the document instance, the fact that the same document is often found in multiple formats (HTML, XML, PDF, etc.), and due to parsing errors which can abruptly skip or truncate portions of the document’s text.
We build a balanced dataset of 24,000 documents for use in tuning hyperparameters for each deduplication algorithm. We further evaluate each technique using its best hyperparameter settings on datasets of 50,000 documents with increasing levels of duplication, from 10% to 90% duplicate instances.
@@ 5.1.5. Choice of values for Tuning Parameters
We evaluate each technique over a range of evenly spaced parameter settings using a grid search. Specifically, for any threshold parameters such as the Jaccard similarity threshold for MinHashLSH and the text overlap threshold for DataComp-LM, we evaluate values between 0.2 and 1.0 in steps of 0.2. We vary the number of permutations for MinHashLSH from 32 to 256 by powers of two. For n-gram techniques, we explore n-gram sizes 1, 2, 5, 7, 13, and 26. For MinHashLSH and LSHBloom, we additionally gathered results for $T=0.5$ and 48 random permutations as the surrounding parameter values demonstrated sharp jumps in the F1 score, and we wanted to explore this change in the score at a finer granularity. All Bloom filters (for DOLMA/DOLMA-ngram, DataComp-LM, LSHBloom) are assigned a false positive rate $p$ of 1e-5. For LSHBloom, this equates to $p\textsubscript{effective}$ between 1e-5 and 8e-4, depending on the value of $T$ and the number of MinHash permutations. The value of $p_{effective}$ at various parameter settings for this benchmark is illustrated in Figure 2.
Figure 2. Value of $p_{effective}$ (the effective false positive rate overhead) in our correctness benchmark at various parameter settings given NN=24,956 documents and pp=1e-5
It is possible that one could properly tune these deduplication methods further using more incremental changes in their parameter settings. However, the expensive runtime of n-gram techniques on large corpora coupled with the need to comb over parameter settings in small increments points to the inherent difficulty and cost associated with properly tuning these methods for maximal deduplication performance. Meanwhile, MinHashLSH-based techniques readily admit parameter tuning with relatively low runtime and memory costs. Furthermore, MinHashLSH and LSHBloom both benefit from established theoretical analysis that effectively bounds their false positive and false negative rates given the values for the Jaccard similarity threshold and number of permutations. MinHashLSH-based techniques offer the greatest opportunity to tune parameters for the best deduplication performance, particularly when faced with a constrained parameter tuning budget. This is particularly true for LSHBloom, which is significantly cheaper to run than MinHashLSH in terms of wall clock time and disk usage—particularly as we increase the size of our dataset to several million or several billion documents (see subsubsection 5.4.1).
(a) MinHashLSH
(b) LSHBloom
Figure 3. F1 score for LSH techniques as a function of the number of permutations (x-axis) and Jaccard similarity threshold (y axis).
5.2. Hyperparameter Tuning Results (24K Dataset)
@@ 5.2.1. LSH Methods
The parametrization of methods such as MinHashLSH and LSHBloom has dramatic effects for their accuracy so tuning the parameters is a critical aspect of deploying them for large-scale deduplication efforts. Figures 3(a) and 3(b) show the F1 score for MinHashLSH and LSHBloom, respectively, with different settings for the Jaccard similarity threshold, $T$, and the number of permutation functions for MinHashing on the 24K dataset. As the number of permutation functions increases, the MinHash estimate of the Jaccard similarity becomes more accurate, and performance generally improves, which is consistent with the theory of MinHashing (broder1998min). Correspondingly, we use 256 MinHash permutations in our subsequent experiments. The threshold parameter affects the F1 score the most, as it defines the sensitivity of the deduplication algorithm to overlap in text content. For our particular benchmark, a threshold of 0.5 yields the best performance; more stringent thresholds lead to a diminishing F1 score.
(a) DOLMA
(b) DataComp-LM
@@ 5.2.2. Paragraph Level and N-Gram Methods
Figure 5 shows the impact of adjusting the overlap threshold parameter for paragraph-level methods, DOLMA and CCNet, on the F1 score. Distinguishing documents at the granularity of individual paragraphs is highly error-prone and results in poor performance. The best F1 for paragraph-level methods is achieved by setting a low threshold of 0.2. N-gram methods such as DCLM do comparably better in terms of F1 score but still lag just behind the scores of the MinHashLSH-based techniques. Interestingly, DOLMA and DOLMA-ngram perform very similarly. Figures 4(a) and 4 show the results of varying both the n-gram size and overlap threshold parameters for each n-gram method. While n-gram methods enable a finer-grained comparison between individual documents and the rest of the corpus, these methods evaluate overlap by determining the portion of n-grams in a document that is exactly duplicated in the corpus. This is similar to the notion of evaluating the set overlap between documents, but can be sensitive to the relative frequency of n-grams within a document, resulting in increased error. Moreover, paragraph-level techniques demonstrate a lack of sensitivity to changes in their parameter settings in this benchmark, indicating bias in their predictions. Specifically, we find that paragraph-level methods tend to be biased toward marking documents as duplicates.
@@
Figure 5. F1 score vs. threshold for paragraph-level techniques.
@@
5.3. Deduplication Correctness (50K Dataset)
(a) Precision
(b) Recall
(c) F1
Figure 6. Correctness metrics for deduplication techniques for various levels of duplication.
@@ 5.3.1. Deduplication Quality
Here we evaluate the correctness of each technique using metrics such as their precision, recall, and F1 score. We do so by constructing datasets of 50,000 documents each with varying levels of duplication between 10% and 90%. We construct these datasets using the procedure described in subsubsection 5.1.4.
Figure 6 illustrates the precision, recall, and F1 score of each technique as we vary the proportion of duplicate documents in our dataset. Notably, in terms of F1 score (see 6(c)), both MinHashLSH and LSHBloom outperform all other methods—only being slightly outperformed by DCLM and Dolma-Ngram (whose performance tracks closely together) when the vast majority ($60\%$) of the dataset consists of duplicates, which is highly unlikely in real-world scenarios. MinHashLSH and LSHBloom also achieve nearly identical F1 scores across experimental settings, exemplifying that our technique can closely match the performance of MinHashLSH while affording considerable gains in scalability (to be shown below in subsection 5.4).
Looking more closely, 6(a) illustrates that MinHashLSH and LSHBloom tend to consistently outperform the other techniques in terms of precision, which is particularly noteworthy for those settings where duplicates are rare, such as when duplicates only comprise 10 or 20% of the dataset. The next best technique in terms of precision is Dolma’s paragraph-level deduplication technique, with its precision often tracking closely with that of the MinHash-based techniques. However, Dolma’s dismal recall scores (<0.4<0.4) suggest that this is likely the result of under-identifying duplicate instances rather than being a reflection of greater capacity for discernment. DCLM, Dolma-Ngram, and CCNet tend to underperform the other techniques in terms of precision. Note that there is a natural tendency for precision scores to increase as the proportion of duplicates in the dataset increases. This is due to a greater number of true positives being present in the dataset.
In terms of recall, 6(b) demonstrates that DCLM and Dolma-Ngram tend to perform the best—though notably their relatively stable recall scores come at the cost of decreased precision, particularly when duplicates are rare. This indicates that, while these techniques identify most duplicates, they struggle to differentiate between duplicate and non-duplicate entities. Meanwhile, MinHashLSH and LSHBloom provide virtually identical recall scores to one another between 0.7 and 0.9 while simultaneously demonstrating strong precision. Paragraph-level techniques demonstrate lower recall than all other techniques by a wide margin.
@@
5.3.2. Performance vs. Quality Tradeoffs
Figures 7(a) and 7(b) plot the F1 score for each deduplication technique against runtime and disk usage, respectively. MinHashLSH and LSHBloom dominate the alternative techniques, achieving the highest F1 scores while also exhibiting competitive runtime. LSHBloom runs faster than MinHashLSH and this runtime advantage becomes more exaggerated at scale (see subsection 5.4). LSHBloom also displays a comparable F1 score to traditional MinHashLSH while using just a fraction of the disk space for its index. As shown below in subsection 5.4, this is a crucial distinction, as rapidly increasing disk usage renders deduplication efforts infeasible at the scale of many millions or billions of documents.
Our results indicate that LSHBloom intelligently discerns between duplicate and non-duplicate documents (as demonstrated by its high precision scores) and can adeptly identify the majority of duplicate documents (as demonstrated by its strong recall scores). Our results also illustrate that the main difference in performance between MinHashLSH and LSHBloom comes in the form of slightly reduced precision for the latter. Intuitively, this makes sense as LSHBloom yields slightly more false positive predictions due to its use of the Bloom filter. Overall, however, both yield similar F1 scores. This, combined with LSHBloom’s greatly reduced resource usage compared to MinHashLSH, positions it as the best choice for text deduplication use-cases.
5.4. Deduplication Resource Usage at Scale
@@
As LLM training datasets grow ever larger, it is important that text deduplication strategies can scale effectively to larger quantities of documents. Here we work with the peS2o dataset (peS2o) of roughly 39M documents. We evaluate each deduplication method on subsets of peS2o of various sizes and compare them in terms of wall clock time and disk usage (e.g., for storing an index or Bloom filter). We do not evaluate the n-gram techniques for scaling capacity on peS2o due to their prohibitive cost. DataComp-LM, for example, requires almost three hours just to deduplicate 50,000 examples in our previous benchmark in subsection 5.3. Even processing just 10% of peS2o using this technique would require over 200 hours by extrapolation. We discuss the results for other methods below.
Figure 8. Disk size of index for different methods, as measured for peS2o subsets of increasing size.
Figure 9. Wall clock times for different methods, as measured for peS2o subsets of increasing size.

### v4

@@ 5. Evaluation
In evaluating a deduplication method for a particular application, we are concerned with its deduplication fidelity, throughput, and disk space requirements. We measure deduplication fidelity via the precision and recall metrics. Low precision, i.e., excessive false positives, is harmful as it unnecessarily reduces dataset size and, in deep learning, limits an LLM’s ability to learn the underlying distribution of the text. Low recall, i.e., excessive false negatives, admits a great degree of duplicate content, with negative effects on LLM quality in pre-training applications. High throughput and a small disk footprint is critical due to the extreme size of modern LLM datasets. For example, peS2o (Soldaini and Lo, 2023) contains roughly 39 million documents, DOLMA (Soldaini and others, 2024) 5B documents, SmolLM3’s pretraining dataset has over 15B documents (Team, 2025), and RedPajama (Computer, 2023) has over 100B documents—scales at which naive pairwise comparisons are infeasible. Therefore, deduplicating at the billion-record scale with high fidelity is paramount.
To assess deduplication fidelity, we tune LSHBloom and other state-of-the-art deduplication methods on a synthetic, balanced dataset of 24K PDFs containing labeled duplicates, we refer to this as our tuning dataset. We then use the best parameter settings for each algorithm and compare them on a larger synthetic dataset of 50K PDFs which we term the testing dataset. Without sufficient fidelity, the system risks either under-identifying duplicates, leading to decreased downstream model quality and unnecessary processing overhead, or over-identifying duplicates, which unnecessarily reduces training dataset volumes. To measure performance, we consider the storage utilization and runtime of each method on a larger collection of 39M PDFs, peS2o (Soldaini and Lo, 2023). Without a sufficiently high throughput and sufficiently low disk footprint, it may not be possible to process large datasets on available hardware resources.
In the following subsections, we describe the details of our evaluation, beginning with the evaluation methodology, then proceeding to our evaluation of fidelity, performance, and finally the tradeoffs between performance and fidelity.
@@ 5.1. Experimental Methodology
We describe our experimental methodology, including testbed hardware, state-of-the-art methods against which we compare, evaluation metrics, and datasets.
@@ 5.1.2. Baselines
We compare LSHBloom with state-of-the-art deduplication methods widely used in large-scale LLM training: MinHashLSH (Broder, 1997), Dolma and Dolma-Ngram (Soldaini and others, 2024), CCNet (Wenzek et al., 2020), and DataComp-LM (Li et al., 2024) (see 3.3 for details). We normalized all implementations to Python implementations to avoid performance differences for common procedures due to implementation language. This also enabled us to normalize the Bloom filter modules for those techniques that use them to a fast C/Python-based implementation provided by pybloomfiltermmap3 (Sinha, 2025). LSHBloom is evaluated using the optimized hashing routine described in Section 4.4. Below, we describe any changes or configurations to the methods.
MinhashLSH (Broder, 1997): We configure traditional MinhashLSH as described in Section 2.3. We use the implementation offered by the popular datasketch package (Zhu, 2025), which implements a hashmap-based LSHIndex. We compute MinHashes using multiprocessing for both MinHashLSH and LSHBloom. DOLMA: We apply both standard Dolma and Dolma-Ngram methods to document-level deduplication. For Dolma, we determine whether a document is duplicated by identifying duplicated paragraphs and comparing the percentage of document text duplicated to the overlap threshold, $T$. For Dolma-Ngram, we split each document into n-grams, determine how many n-grams are duplicated according to the Bloom filter, and compare the percentage of n-grams duplicated within that document to the overlap threshold, $T$. Our implementation aims to faithfully extend Dolma’s two primary deduplication methods to document-level deduplication. CCNet originally performs deduplication at the paragraph level; we apply it at the document level by measuring the proportion of duplicated paragraphs in any given document against some tolerance threshold, $T\in$ [0, 1]. We augment CCNet with multiprocessing where appropriate (e.g., when computing hashes over each document) for a fair comparison during our scaling tests. DataComp-LM: For our evaluation, we compare only with their document-level deduplication procedure.
For Dolma, Dolma-Ngram, and DCLM, we must give an expected n-gram/paragraph count for the corpus; choosing this count accurately is crucial to making a fair comparison because it influences Bloom filter size and error rates. Because counting the number of n-grams/paragraphs in a corpus is very expensive compared to counting documents (as is required for MinHash-based techniques), we compute an estimate of the n-gram/paragraph count in a corpus by randomly sampling $N$ = 1000 documents from the corpus, computing the mean n-gram/paragraph count from our sample, and multiplying the mean count by the number of total documents in that corpus to obtain an estimated total n-gram/paragraph count.
@@ 5.1.3. Metrics
Fidelity: We require methods that can distinguish meaningfully between duplicate and non-duplicate documents. Here, we consider a ”positive” label to mean that a document is a duplicate of another document in the corpus. We want these methods to yield few false positives (to retain the maximal number of informative records), few false negatives (to reduce the prevalence of duplication), and to provide a reasonable combination of accuracy and performance. Particularly for large-scale deep learning, retaining the maximal number of training examples is highly desirable—therefore, we want few false positives. Similarly, allowing a large number of duplicates into the training corpus is unacceptable as this affects the model’s training efficiency and decreases model quality. To account for these considerations, we evaluate each deduplication technique based on its ability to mitigate both false positives and false negatives using its precision (the proportion of duplicate labels that are true duplicates), its recall (the proportion of duplicates in the dataset that we correctly identify), and its F1 score (the harmonic mean of precision and recall):
@@
Performance: To assess the resource costs of each technique, we measure wall clock time (using /user/bin/time) and disk space requirements (using Python’s standard os.path.getsize).
5.1.4. Constructing the fidelity evaluation dataset
To permit the rigorous evaluation of de-duplication methods, we created a pipeline for generating labeled datasets of duplicate documents. Siebenschuh et al. (2025) produced a dataset of parsed PDFs for evaluating their adaptive PDF parsing technique, AdaParse. We utilize this baseline dataset to construct synthetic tuning and testing datasets for the deduplication task. The baseline dataset consists of scientific articles with two versions: one scraped from HTML and the other parsed from PDF. These span diverse disciplines (mathematics, physics, chemistry, biology, economics, engineering, and medicine) as identified through keywords or classification algorithms based on metadata. Text content from HTML is directly extracted. Each PDF is parsed by applying a text extraction or OCR tool—one of PyMupDF (Smith, 2024), Nougat (Blecher et al., 2023), or Tesseract (Smith, 2007), each used with roughly the same frequency. We construct datasets containing duplicated documents from this collection of parsed texts, either by including multiple versions of the same document parsed with different tools and/or by including randomly truncated versions of that document. We balance the number of parser and truncation-based duplicates in each dataset we construct, which prevents evaluation bias towards techniques that are better suited to identifying just one of these types of duplication. These scenarios reflect common forms of duplication encountered in real-world web scraping pipelines where the same document may be ingested via different text extraction tooling due to differences in document origin, the same document being obtained in multiple formats (HTML, XML, PDF, etc.), and due to parsing errors, which can abruptly skip or truncate portions of the document’s text.
We build a balanced dataset of 24,000 documents for use in tuning hyperparameters for each deduplication algorithm, the tuning dataset. We further evaluate each technique using its best hyperparameter settings on datasets of 50,000 documents with increasing levels of duplication, from 10% to 90% duplicate instances, the testing datasets.
@@ 5.1.5. Choice of values for Tuning Parameters
We evaluate each technique for a range of parameter settings. Specifically, for each threshold parameter, such as the Jaccard similarity threshold for MinHashLSH and the text overlap threshold for DataComp-LM, we evaluate values from 0.2 to 1.0 in steps of 0.2. We vary the number of permutations for MinHashing from 32 to 256 by powers of two. For n-gram techniques, we explore n-gram sizes 1, 2, 5, 7, 13, and 26. For MinHashLSH and LSHBloom, we also gathered results for $T$ = 0.5 and 48 random permutations, as the surrounding parameter values demonstrated sharp jumps in F1 score, and we wanted to explore this change in the score at a finer granularity. All Bloom filters (for DOLMA/DOLMA-ngram, DataComp-LM) are assigned a false positive rate $p$ of 1e-5. For LSHBloom, we set $p\textsubscript{effective}$ to 1e-5.
It is possible that one could properly tune these deduplication methods further using more incremental changes in their parameter settings. However, the expensive runtime of n-gram techniques on large corpora, coupled with the need to comb over parameter settings in small increments points to the inherent difficulty and cost associated with properly tuning these methods for maximal deduplication performance. Meanwhile, MinHashLSH-based techniques readily admit parameter tuning with relatively low runtime and memory costs. Furthermore, MinHashLSH and LSHBloom both benefit from established theoretical analysis that effectively bounds their false positive and false negative rates given the values for the Jaccard similarity threshold and number of permutations (Leskovec et al., 2014). MinHashLSH-based techniques offer the greatest opportunity to tune parameters for the best deduplication performance, particularly when faced with a constrained parameter tuning budget. This is particularly true for LSHBloom, which is significantly cheaper to run than MinHashLSH in terms of wall clock time and disk usage—particularly as we increase the size of our dataset to several million or several billion documents (see Section 5.4.1).
Figure 2. F1 score for LSH techniques as a function of number of permutations (x-axis) and Jaccard similarity threshold (y axis).
5.2. Hyperparameter Tuning
@@ 5.2.1. LSH Methods
The parametrization of methods such as MinHashLSH and LSHBloom has dramatic effects on their accuracy, so tuning the parameters is a critical aspect of deploying them for large-scale deduplication efforts. Figure 2 shows the F1 score for MinHashLSH and LSHBloom, respectively, with different settings for the Jaccard similarity threshold, $T$, and the number of permutation functions for MinHashing on the tuning dataset. As the number of permutation functions increases, the MinHash estimate of the Jaccard similarity becomes more accurate, and performance generally improves, which is consistent with the theory of MinHashing (Broder et al., 1998). Correspondingly, we use 256 MinHash permutations in our subsequent experiments. The threshold parameter affects the F1 score the most, as it defines the sensitivity of the deduplication algorithm to overlap in text content. For our particular benchmark, a threshold of 0.5 yields the best performance; more stringent thresholds lead to a diminishing F1 score.
@@ 5.2.2. Paragraph Level and N-Gram Methods
Figure 3 shows the results of varying both the n-gram size and overlap threshold parameters for each n-gram method. DCLM is fairly successful, even approaching the F1 score of the MinHashLSH-based techniques, and demonstrates moderate sensitivity to the overlap threshold parameter. Smaller n-gram sizes seem to perform better as well—likely due to the increased granularity with which they enable comparison between a document and the remainder of the corpus. For DCLM, the best setting is using an overlap threshold of 0.2 and an n-gram size of 5. In contrast, Dolma-Ngram seems to be fairly insensitive to the choice of parameters and performs relatively poorly. The difference between these two very similar techniques may stem from DCLM’s use of the UniSeg tokenizer, which splits text into meaningful units based on rules defined by the Unicode standard—Dolma-Ngram simply splits text by whitespace.
Figure 4 shows the impact of adjusting the overlap threshold parameter for paragraph-level methods, DOLMA and CCNet, on F1 score. Predictably, distinguishing documents at the granularity of individual paragraphs is highly error-prone and results in poor performance. The best F1 for paragraph-level methods is achieved by setting a low threshold of 0.2. While n-gram methods enable finer-grained comparison between individual documents and the rest of the corpus, these methods evaluate overlap by determining the portion of n-grams in a document that is exactly duplicated in the corpus. This is somewhat similar to the notion of evaluating the set overlap between documents, but can be sensitive to the relative frequency of n-grams within a document, resulting in increased error. Moreover, paragraph-level techniques demonstrate a lack of sensitivity to changes in their parameter settings in this benchmark, indicating bias in their predictions. Specifically, we find that paragraph-level methods tend to be biased toward marking documents as duplicates.
@@
Figure 4. F1 score vs. threshold, paragraph-level techniques.
@@
5.3. Deduplication Fidelity
Figure 5. Fidelity metrics for deduplication techniques for various levels of duplication.
@@ 5.3.1. Deduplication Quality
Here we evaluate the fidelity of each technique using metrics such as precision, recall, and F1 score. We do so by constructing datasets of 50,000 documents each with varying levels of duplication between 10% and 90%. We construct these datasets using the procedure described in Section 5.1.4.
Figure 5 illustrates the precision, recall, and F1 score of each technique as we vary the proportion of duplicate documents in our dataset. Notably, in terms of F1 score, both MinHashLSH and LSHBloom outperform all other methods—only slightly outperformed by DCLM and Dolma-Ngram (whose performance tracks closely together) when the vast majority (>>60%) of the dataset consists of duplicates, highly unlikely in real-world scenarios. MinHashLSH and LSHBloom also achieve nearly identical F1 scores across experimental settings, exemplifying that our technique can closely match MinHashLSH’s performance while affording considerable gains in scalability (to be shown below in Section 5.4).
Looking more closely, MinHashLSH and LSHBloom tend to consistently outperform the other techniques in terms of precision, which is particularly noteworthy for those settings where duplicates are rare, such as when duplicates only comprise 10 or 20% of the dataset. The next best technique in terms of precision is Dolma’s paragraph-level deduplication technique, with its precision tracking closely with that of the MinHash-based techniques. However, Dolma’s dismal recall scores (<0.4<0.4) suggest that this is likely the result of under-identifying duplicate instances rather than being a reflection of greater capacity for discernment. DCLM, Dolma-Ngram, and CCNet tend to underperform the other techniques in terms of precision. Note that there is a natural tendency for precision scores to increase as the proportion of duplicates in the dataset increases. This is due to a greater number of true positives being present in the dataset.
In terms of recall, DCLM and Dolma-Ngram tend to perform the best, although their relatively stable recall scores come at the cost of decreased precision, particularly when duplicates are rare. This indicates that, while these techniques identify most duplicates, they struggle to differentiate between duplicate and non-duplicate entities. Meanwhile, MinHashLSH and LSHBloom provide virtually identical recall scores to one another between 0.7 and 0.9 while simultaneously demonstrating strong precision. Paragraph-level techniques demonstrate lower recall than all other techniques by a wide margin.
@@
5.3.2. Performance vs. Quality Tradeoffs
Figures 6(a) and 6(b) plot the F1 score for each deduplication technique against runtime and disk usage, respectively. MinHashLSH and LSHBloom dominate the alternative techniques, achieving the highest F1 scores while also exhibiting competitive runtime. LSHBloom runs faster than MinHashLSH and this runtime advantage becomes more exaggerated at scale (see Section 5.4). LSHBloom also displays a comparable F1 score to traditional MinHashLSH while using just a fraction of the disk space for its index. As shown below in Section 5.4, this is a crucial distinction, as rapidly increasing disk usage renders deduplication efforts infeasible at the scale of many millions or billions of documents.
Our results indicate that LSHBloom intelligently discerns between duplicate and non-duplicate documents (as demonstrated by its high precision scores) and can adeptly identify the majority of duplicate documents (as demonstrated by its strong recall scores). Our results also illustrate that the main difference in performance between MinHashLSH and LSHBloom comes in the form of slightly reduced precision for the latter. Intuitively, this makes sense as LSHBloom yields slightly more false positive predictions due to its use of the Bloom filter. Overall, however, both yield similar F1 scores. This, combined with LSHBloom’s greatly reduced resource usage compared to MinHashLSH, positions it as the best choice for text deduplication.
5.4. Deduplication Resource Usage at Scale
As LLM training datasets grow ever larger, it is important that text deduplication strategies can scale effectively to larger quantities of documents. Here we work with the peS2o dataset (Soldaini and Lo, 2023) of roughly 39M documents. We evaluate each deduplication method on subsets of peS2o of various sizes and compare them in terms of wall clock time and disk usage (e.g., for storing an index or Bloom filter). We do not evaluate the n-gram techniques for scaling capacity on peS2o due to their prohibitive cost. DataComp-LM, for example, requires almost three hours just to deduplicate 50,000 examples in our previous benchmark in Section 5.3. Even processing just 10% of peS2o using this technique would require over 200 hours by extrapolation. We find that Dolma-Ngram is similarly prohibitive in runtime, likely due to the time cost of processing many n-grams in documents that are frequently much larger than those in our previous benchmark. We discuss the results for other methods below.
(a) Wall clock times over subsets of peS2o.
(b) Disk Usage over subsets of peS2o.
Figure 7. Resource usage at scale for each deduplication technique.
Figure 8. Extrapolated wall clock times per method at scale.